# Generative AI and Firm Productivity:

### v2

@@ 5.3 Heterogeneous Effects Across Products
The platform hosts hundreds of millions of products spanning diverse categories. To study heterogeneity across products, we classify them into high and low groups based on three pre-experiment metrics. First, category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. This metric reflects the extent of product and demand differentiation, with higher concentration indicating more standardized offerings and relatively homogeneous consumer preferences (e.g., laptops typically exhibit higher concentration than dresses, where product features and consumer tastes are more varied). Products in the low group belong to categories with concentration below the platform average. Second, annual sales quantity, defined within category. Products in the low group—referred to as tail products—comprise the bottom 50% cumulative share of units sold when ranked by annual sales quantity within a category. Third, product price, also defined within category to account for category-specific pricing variation. Low-priced products are those priced below the median of their respective category. Summary results are reported in Table 9, with additional details in Appendix F.
Table 9: Summary of Heterogeneous Treatment Effects Across Products
(1) (2) (3) (4) (5) (6)
Business Dependent Market Concentration Annual Quantity Price
Workflow Variable High Low High Low High Low
1 Pre-sale Sales 21.3%* 13.2%* 8.8% 21.3%*** 16.6%* 15.9%*
Serv Chatbot Conversion 26.6%*** 18.5%*** 10.3% 30.3%*** 17.4%* 24.5%***
2 Search Query Sales -0.80% 6.49%*** -0.47% 4.92%*** 3.79%** 0.89%
Refinement Conversion 0.45% 1.85%*** 0.00% 2.07%*** 1.19%* 0.94%
3 Product Sales 3.10%*** 0.55% 2.10%* 2.03%* 4.10%*** 0.89%
Description Conversion 1.38%** 0.81% 2.07%*** 0.69% 3.32%*** 0.63%
4 Marketing Sales -0.8% 5.8% -0.9% 2.4% -5.2% 4.9%
Push Message Conversion 2.9% 3.0%* 0.8% 7.0%*** 4.1%*** 2.7%
5 Google Sales -4.3% -5.2% -2.3% -8.9% -6.2% -2.3%
Ad Titles Conversion -4.1% -2.1% -2.9% -4.4% -4.2% -2.8%
  • 1
    We classify products into high and low groups based on three pre-experiment metrics. (1) Category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. Products in the low group belong to categories with concentration levels below the platform average. (2) Annual sales quantity, also defined within category. Products in the low group—referred to as tail products—are those comprising the bottom 50% cumulative share of total units sold when ranking products by annual sales quantity within each category. (3) Product price, defined within each category to control for category-level pricing variation. Low-priced products are those priced below the median of their respective category.
  • 2
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
  • 3
    We report the % change. % change is calculated by dividing the treatment effect by the control group average. The asterisks notation indicates the significance of treatment effect. *** p<<0.01, ** p<<0.05, * p<<0.1.
Pre-sale Service Chatbot.
GenAI-enabled pre-sale service chatbot exhibits larger effects on sales and conversion rates for tail products with low sales volume (Column 4). A plausible explanation is that such products often suffer from poor visibility and information gaps that heighten consumer hesitation [32]. GenAI chatbot mitigates these frictions by clarifying product details and providing richer information before the purchase, particularly for niche or long-tail items with weak marketing signals. Heterogeneity across prices is more nuanced, with the intervention benefiting both high- and low-priced products (Columns 5 and 6). For high-priced items, the chatbot likely supports conversion by helping consumers justify larger expenditures through detailed explanations of value, quality assurance, and comparative advantages. For low-priced goods, the chatbot reduces decision frictions by streamlining information provision and simplifying the purchase process. Similarly, heterogeneity across categories defined by market concentration is also nuanced (Columns 1 and 2), indicating that the GenAI-augmented chatbot benefits products in both highly concentrated and less concentrated categories.
Search Query Refinement.
For this experiment, our results show significant increases in sales and conversion rates only for products in low-concentration categories (Column 2), tail products (Column 4), and high-priced segments (Column 5). These patterns are consistent with the idea that GenAI reduces search friction by enhancing the comprehension and articulation of consumer demand in query-based search engines. This effect is particularly valuable in categories with high product and demand differentiation, because in such settings, consumers often struggle to express their needs precisely, yet accurate and content-clear queries are critical for retrieving well-matched products. For example, searching for a laptop is relatively straightforward, as the category is concentrated around key brands and standardized specifications, whereas searching for a dress is more complex given the wide variation in designer, color, texture, style, and other subjective attributes. In addition, the reduction in frictions can also increase the visibility of long-tail products and strengthening purchase determination of high-priced items.
Product Description.
We find significant gains from GenAI-generated product descriptions in high-concentration categories (Column 1), in high-priced products (Column 5), and across both head and tail products (Columns 3 and 4). These patterns are consistent with several mechanisms. In high-concentration categories such as laptops, product attributes are more standardized and therefore easier to communicate through text-based descriptions. For high-priced items, more comprehensive descriptions may help mitigate information asymmetry, which is particularly salient in costly purchase decisions. The results for head versus tail products may partly reflect the experimental context: the intervention was implemented on platform self-sold products, where Chinese vendors typically provide image-based content that often misaligns with cross-border consumer preferences. Consequently, the head–tail classification may not accurately reflect the quality of text-based description content. For instance, a head product may feature high-quality image-based content but still lack a well-developed text description.
Marketing Push Message.
Our results show no significant effects of GenAI deployment on sales across product segments, consistent with the main analysis. However, we find significant conversion gains in categories with low market concentration (Column 2), in tail products (Column 4), and in high-priced products (Column 5). These patterns align with intuitive mechanisms. In categories such as apparel and fashion, where products are highly differentiated and consumer choices hinge on subtle differences in style, quality, and branding, GenAI-powered marketing—especially when personalized—can more effectively capture attention. Tail products, which typically lack historical data or reputational signals, also benefit more from enhanced marketing outreach. For high-priced items, tailored messaging helps emphasize value, quality assurance, and differentiation, making premium offerings more persuasive.
Google Advertising Title.
All heterogeneity results are statistically insignificant in this experiment.
Overall, our analysis of product heterogeneity shows that the effects of GenAI applications are context-dependent. Search Query Refinement and Marketing Push Message generate larger improvements in low-concentration categories, where product differentiation is greater and consumer preferences more diverse. By contrast, Pre-Sale Service Chatbot and Product Description yield stronger gains in highly concentrated categories. Segmenting products by annual sales and price further reveals more consistent patterns: across most workflows, GenAI tends to deliver larger benefits for tail products and for high-priced items.
@@ 6 Discussion and Conclusions
Our findings yield three main insights. First, GenAI can deliver measurable productivity improvements with comparable inputs, as reflected in increased sales across several business workflows. While the magnitude of these gains varies widely—from negligible effects to double-digit increases—the evidence demonstrates that GenAI can generate substantial economic value when deployed in consumer-facing processes. Second, these improvements arise primarily from enhanced consumer experience and satisfaction through the GenAI-driven reduction of frictions in the marketplace, rather than from cost savings on the input side. Across workflows, we observe higher conversion rates: by enriching pre-sale communication, refining search queries, generating richer product descriptions, and personalizing marketing messages, GenAI improves matching efficiency and mitigates information asymmetries. Third, the benefits of GenAI adoption are heterogeneous. Smaller sellers, less experienced consumers, tail products, and high-priced items derive disproportionately larger gains, highlighting GenAI’s role in bridging capability gaps across different segments of the marketplace.
@@
Because most experiments were randomized at the consumer level and overlap across experiments was minimal (less than 1%), the observed effects capture incremental demand (i.e., market expansion) rather than substitution across products. Back-of-the-envelope calculations—annualizing workflow-specific gains and assuming linear additivity—suggest that the four GenAI applications with positive sales effects generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects represent roughly 5.5–6% of global per-user e-commerce revenue growth in 2023–2024. Taken together, these figures show that even a few GenAI applications already yield substantial gains for a large, mature retailer, with the potential for much larger effects as adoption broadens and increasingly targets revenue-critical workflows. By 2025, the partner platform had deployed GenAI in more than 60 workflows, with usage rising twentyfold as API calls to its proprietary GenAI tools increased between 2024 and 2025.
@@
At the same time, our study has several limitations that inform the interpretation of the results and point to avenues for future research. First, the adoption horizon in our experiments was short—spanning several weeks to months—so we capture only the immediate, short-run effects of GenAI. Impacts of sustained GenAI use may differ as sellers and consumers adapt their behavior and as platforms refine model deployment. Relatedly, we lack data to assess long-term consumer responses, such as product returns and retention. For instance, while GenAI may initially boost purchases, consumers may ultimately become dissatisfied if the AI-generated content fails to align with actual product characteristics. Second, our analysis is limited to seven workflows that were selected by the platform based on managerial assessments of technical feasibility, organizational costs, and expected productivity improvements, rather than representing the full spectrum of business processes where the technology could be deployed. Other business processes, including logistics, inventory management, or dynamic pricing, remain unexplored and could yield distinct productivity effects. Third, while our estimates map directly into total factor productivity gains under the assumption of constant inputs, we cannot rule out future changes in labor and capital inputs. Many of the studied processes—such as customer service, content creation, and Chargeback Defense—are currently staffed or supported by human labor. Over time, GenAI adoption could displace or augment these functions, yielding additional cost-side efficiency improvements not captured in our current analysis.
@@
Another limitation concerns external validity in general equilibrium. Our experiments were conducted on a single, albeit large, global retail platform. The effects we document partly reflect relative improvements in user experience and satisfaction within this environment. If GenAI adoption becomes widespread and competing platforms deploy similar technologies, these relative advantages may diminish. However, if the mechanism we highlight—enhanced consumer experience through reduced frictions—holds more broadly, there may still be scope for market expansion even in a more competitive environment. In addition, our experiments abstract from strategic responses by competitors —such as changes in pricing or advertising strategy— which could either amplify or attenuate realized productivity gains.
@@
Taken together, our results highlight both the current scope and the future potential of GenAI adoption in online retail. In the short run, GenAI delivers measurable productivity gains both within workflows and across the platform by reducing market frictions and enhancing the consumer shopping experience, constituting a substantial contribution given the scale and maturity of our focal platform. Over the longer run, its transformative impact could grow if firms expand adoption beyond early-use cases, capture cost-reduction opportunities, and adapt organizational structures to integrate GenAI more effectively. Continued advances in computational speed, accuracy, and domain coverage would further amplify this potential. At the aggregate level, widespread industry adoption raises open questions about equilibrium effects, competitive dynamics, and the durability of observed gains. Our study offers a first step by providing causal evidence on how GenAI can reengineer core retail workflows to deliver meaningful business outcomes, while pointing to important avenues for future research on general equilibrium impacts, cost-side adjustments, and long-term productivity growth.
@@ Appendix A Illustrative User Interfaces and Examples of Each Experiment
Pre-sale Service Chatbot.
@@
Search Query Refinement.
@@
Product Description.
In Figure A3, Panel (a) illustrates the control condition with human-generated descriptions, while Panel (b) shows the treatment condition, where GenAI-created descriptions are layered on top of those written by humans. The GenAI-generated content provides more comprehensive and structured bullet-point-style descriptions that highlight product features, benefits, and typical use cases.
@@
Marketing Push Message.
@@
Google Advertising Title.
@@
Live Chat Translation.
@@ Appendix B Covariance Balance Checks for the Field Experiments
    Mean (Std. Dev.) are shown with all values normalized by the corresponding variable’s control group mean. For the definitions of each variable, please refer to the notes in Figure 1.
@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
Pre-sale Service Chatbot.
Pre-sale inquiries regarding product and seller information (e.g., product attributes, promotions, and logistics) play a critical role in shaping consumer purchase decisions. To support decision-making, improve consumer service, and enhance the overall consumer experience, the platform introduced a GenAI-powered chatbot capable of delivering accurate, content-rich responses tailored to a diverse consumer base and available around the clock.
@@
In Table C1, Columns (1)–(4) report effects when the treatment is GenAI Reply, and Columns (5)–(8) report effects for GenAI+Human Reply. Within each set, Columns (1)–(2) and (5)–(6) use No Serice as the control, while Columns (3)–(4) and (7)–(8) use Human Reply as the control. Focusing on the comparison between the No Service control and the GenAI Reply treatment (Cols. 1–2), we continue to find sizable productivity improvements: the conversion rate rises by 21.7% and sales increase by 16.3% (both significant at the 1% level). When the No Service control is compared with the GenAI+Human Reply treatment (Cols. 5–6), the gains are even larger—conversion improves by 29.0% and sales by 25.0%—indicating complementarities between GenAI and human agents. Using Human Reply as the control, the GenAI Reply treatment shows no statistically significant differences in either conversion or sales (Cols. 3–4), suggesting that the GenAI chatbot matches the quality of human service but does not outperform it. By contrast, relative to the Human Reply control, the GenAI+Human Reply treatment yields a small, statistically insignificant increase in conversion (4.8%) and a marginally significant 11.5% increase in sales (Cols. 7–8), implying that the hybrid approach can enhance revenue even when benchmarked against human agents. Table C2 reports the Cart Value in the Pre-sale Service Chatbot experiment. We find that all Cart Values are statistically insignificant.
Table C1: Main Effect of Pre-sale Service Chatbot
Control: No Service Human Reply No Service Human Reply
(1) (2) (3) (4) (5) (6) (7) (8)
Conv Rate Sales Conv Rate Sales Conv Rate Sales Conv Rate Sales
Treat 0.0131*** 0.274*** -0.000768 0.0701 0.0175*** 0.422*** 0.00358 0.218*
(0.00256) (0.0995) (0.00261) (0.0992) (0.00295) (0.115) (0.00301) (0.1145)
%Change 21.7% 16.3% -1.0% 3.7% 29.0% 25.0% 4.8% 11.5%
Observations 44,614 44,614 44,736 44,736 30,345 30,345 30,467 30,467
R-squared 0.001 0.000 0.000 0.000 0.001 0.000 0.000 0.000
  • 1
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
  • 2
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Table C2: Cart Value in Pre-sale Service Chatbot Experiment
@@
%Change -3.1% 6.4% -4.5% 4.8%
Observations 2,092 2,316 3,076 3,300
@@
    The dependent variable is the “Cart Value”, which refers to the expenditure per consumer, conditional on making a purchase. All analyses are restricted to those who make a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement.
@@
The results are summarized in Table C3. Column 1 indicated no significant differences in product views between the two groups. However, treatment group consumers generated 1.10% more clicks (Column 2), spent 2.93% more (Column 4), and were 1.15% more likely to make a purchase (Column 5). These effects were primarily driven by a significant 2.02% rise in click-through rate (Column 6)—the ratio of product clicks to views—suggesting that consumers found the exposed products more appealing and chose to seek additional details after viewing the summarized search results. Furthermore, the click-to-order conversion rate (Column 7)—the ratio of product orders to clicks—remained insignificant, echoing the fact that query refinement influenced only the composition of products retrieved immediately after a query search, not the information displayed on product detail pages. Overall, GenAI-facilitated query refinement enhanced the search algorithm’s ability to satisfy consumers’ demand, resulting in more effective matching and improved consumer purchases.
@@
(1) (2) (3) (4) (5) (6) (7) (8)
Click- Click-to-Order
Conversion Cart through Conversion
View Click Order Sales Rate Value Rate Rate
Treat -0.549 0.0901*** 0.00154 0.0648** 0.00101** 0.376 0.000767*** -0.000088
(0.887) (0.0245) (0.00107) (0.0314) (0.000411) (0.334) (0.000079) (0.000182)
%Change -0.18% 1.10% 0.94% 2.93% 1.15% 1.49% 2.02% -0.35%
Observations 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 163,471 1,849,382 1,508,873
R-squared 0.038 0.041 0.019 0.004 0.030 0.010 0.003 0.014
@@
    “View” stands for the number of product views in the search results pages. “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Click-through Rate” is the ratio of the number of product clicks to the number of product views. It measures the degree of conversion from views to clicks. “Click-to-Order Conversion Rate” stands for the ratio of the number of product orders to the number of product clicks. It measures the degree of conversion from clicks to purchases.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Product Description.
@@
In Column 1 of Table C4, we found no significant differences in product clicks between the control and treatment groups. However, once consumers clicked and accessed the product detail pages where descriptions were displayed, the treatment group consumers tended to place 1.08% more orders and spend 2.05% more on those orders (Columns 2 and 3). This improvement was also evidenced by the 1.27% increase in the conversion rate in Column 4, indicating that GenAI-generated descriptions promoted a higher likelihood of purchasing. Consequently, augmenting human-generated product descriptions with GenAI-generated content effectively motivated consumers’ purchase decisions, leading to higher sales for the platform.

### v3

@@ 5.3 Heterogeneous Effects Across Products
Figure 5: Heterogeneous Treatment Effects on Sales Across Products
Chatbot Refer to caption
Search Refer to caption
Description Refer to caption
Push Refer to caption
Google Refer to caption
Notes: The bars indicate the % change in treatment effects for the high (blue) and low (red) groups. The p-values are from seemingly unrelated estimations testing the equality of % treatment effects across the two groups.
Finally, we study heterogeneity in GenAI’s effects across products by classifying them into high and low groups along three pre-experiment dimensions. First, category market concentration, measured by the sales share of the top 1% of products within each category, captures the degree of product standardization and demand homogeneity. Second, annual sales quantity within category distinguishes head from tail products. Third, product price within category proxies for purchase stakes and potential information asymmetries. For each dimension, products in the low group correspond to less concentrated categories, lower-sales (tail) products, or lower-priced items. [^23] Products are classified into the low group if they (i) belong to categories with concentration below the platform average, (ii) account for the bottom 50% of cumulative sales quantity within their category, or (iii) are priced below the category median. We decompose consumer purchases into transactions involving high- and low-group products and estimate treatment effects separately by group, testing for differences in percentage treatment effects using seemingly unrelated estimation.
Summary results are reported in Figure 5, with regression results and additional details provided in Appendix F. Overall, we find no uniform or sharply defined pattern of product-level heterogeneity across workflows. While several point estimates suggest larger GenAI-driven gains for products in less concentrated categories, tail products or high-priced items—depending on the specific application—most differences between high- and low-group products are statistically insignificant. The direction of the effects is often consistent with GenAI reducing search and information frictions in settings with greater product differentiation, limited sales history or higher decision stakes, but these patterns are not robust across all workflows. Overall, the evidence points to limited and context-dependent product-level heterogeneity in the effects of GenAI.
@@ 6 Discussion and Conclusions
Our findings yield three main insights. First, GenAI can deliver measurable productivity improvements while holding inputs constant, as reflected in increased sales across several business workflows. Although the magnitude of these gains varies widely, ranging from negligible effects to double-digit increases, the evidence shows that GenAI can generate economically meaningful value when deployed in consumer-facing processes. Second, these improvements arise primarily from enhanced consumer experience through GenAI-driven reductions in marketplace frictions, rather than from cost savings on the input side. Across workflows, we observe higher conversion rates but no changes in spending intensity or adverse post-purchase outcomes. By enriching pre-sale communication, refining search queries, generating richer product descriptions, and personalizing marketing messages, GenAI improves matching efficiency and mitigates information asymmetries along the customer journey. Third, the benefits of GenAI adoption are heterogeneous across market participants. On the demand side, less experienced consumers derive disproportionately larger gains. On the supply side, point estimates suggest larger benefits for smaller and newer sellers, though these effects are less precisely estimated. Heterogeneity across products is more context-dependent. Taken together, these patterns highlight GenAI’s potential to improve outcomes for participants with lower baseline capabilities.
@@
Because most experiments were randomized at the consumer level and overlap across experiments was minimal (less than 1%), the observed effects capture incremental demand (i.e., market expansion) rather than substitution across products. [^24] One exception is the Product Description experiment, where observed consumer outcomes pertain only to the products included in the experiment and do not capture behavior toward competing products. Back-of-the-envelope calculations—annualizing workflow-specific gains and assuming linear additivity—suggest that the four GenAI applications with positive sales effects generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects correspond to roughly 5.5–6% of global per-user e-commerce revenue growth in 2023–2024. Taken together, these figures show that even a few GenAI applications already yield substantial gains for a large, mature retailer, with the potential for much larger effects as adoption broadens and increasingly targets revenue-critical workflows. By 2025, the partner platform had deployed GenAI in more than 60 workflows, with usage rising twentyfold as API calls to its proprietary GenAI tools increased between 2024 and 2025.
@@
At the same time, our study has several limitations that inform the interpretation of the results and point to avenues for future research. First, the adoption horizon in our experiments was short, spanning several weeks to months. Therefore, our analysis reflects only the immediate, short-run effects of GenAI. The impacts of sustained GenAI use may differ as sellers and consumers adapt their behavior over time and as platforms refine model deployment. Relatedly, we lack data to assess longer-term outcomes such as consumer retention and repeat purchasing. For example, while our results show no deterioration in post-purchase outcomes, as measured by return rates and customer ratings, they do not speak to longer-run consumer responses, such as changes in trust, loyalty, or engagement with the platform. Second, our analysis is limited to seven workflows that were selected by the platform based on managerial assessments of technical feasibility, organizational costs, and expected productivity improvements, rather than representing the full spectrum of business processes where the technology could be deployed. Other business processes, including logistics, inventory management, or dynamic pricing, remain unexplored and could yield distinct productivity effects. Third, while our estimates map directly into total factor productivity gains under the assumption of constant inputs, we cannot rule out future changes in labor and capital inputs. Many of the studied processes, including customer service, content creation, and chargeback defense, are currently staffed or supported by human labor. Over time, GenAI adoption could displace or augment these functions, yielding additional labor efficiency improvements not captured in our current analysis.
@@
Another limitation concerns external validity in general equilibrium. Our experiments were conducted on a single, albeit large, global retail platform. The effects we document partly reflect relative improvements in user experience within this environment. As GenAI adoption spreads and competing platforms deploy similar technologies, these relative advantages may diminish. However, if the mechanism we highlight—higher willingness to purchase due to enhanced consumer experience through reduced frictions—holds more broadly, there may still be scope for market expansion even in a more competitive environment. In addition, our experiments abstract from strategic responses by competitors, such as changes in pricing or advertising strategy, which could either amplify or attenuate realized productivity gains.
@@
Taken together, our results highlight both the current scope and the future potential of GenAI adoption in online retail. In the short run, GenAI delivers measurable productivity gains by creating demand-side value, both within individual workflows and across the platform, through reductions in market frictions and improvements in the consumer shopping experience. These gains are economically meaningful given the scale and maturity of the partner platform. Over the longer run, GenAI’s impact may expand as firms move beyond early use cases, capture cost-reduction opportunities, and adapt organizational structures to integrate the technology more effectively. Continued advances in computational speed, accuracy, and domain coverage are likely to further amplify these effects. At the aggregate level, however, widespread adoption raises open questions about equilibrium effects, competitive dynamics, and the persistence of observed gains. Our study provides a first step by offering causal evidence on how GenAI can reshape core retail workflows to improve meaningful business outcomes, while highlighting important directions for future research on general equilibrium effects, cost-side adjustments, and long-term productivity growth.
@@ Appendix A Illustrative User Interfaces and Examples of Each Experiment
Pre-sale Service Chatbot
@@
Search Query Refinement
@@
Product Description
In Figure A3, Panel (a) illustrates the control condition with human-generated descriptions, while Panel (b) shows the treatment condition, where GenAI-created descriptions are layered on top of those written by humans. The AI-generated content provides more comprehensive and structured bullet-point-style descriptions that highlight product features, benefits, and typical use cases.
@@
Marketing Push Message
@@
Google Advertising Title
@@
Live Chat Translation
@@ Appendix B Covariance Balance Checks for the Field Experiments
    Mean (Std. Dev.) are shown with all values normalized by the corresponding variable’s control group mean. For the definitions of each variable, refer to the notes in Figure 1.
@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
Pre-sale Service Chatbot
Pre-sale inquiries regarding product and seller information (e.g., product attributes, promotions, and logistics) play a critical role in shaping consumer purchase decisions. To reduce search costs, support decision-making, improve customer service, and enhance the overall consumer experience, the platform introduced a GenAI-powered chatbot capable of delivering accurate, content-rich responses tailored to a diverse consumer base and available around the clock.
@@
Table C1 presents the impact of the GenAI chatbot on sales, while Table C2 reports its effects on conversion rates and cart value. In both tables, Columns (1)–(2) report effects when the treatment is GenAI Reply, and Columns (3)–(4) report effects for GenAI+Human Reply. Within each set, Columns (1) and (3) use No Service as the control, while Columns (2) and (4) use Human Reply as the control.
Focusing on the comparison between the No Service control and the GenAI Reply treatment (Column 1), we continue to find sizable productivity improvements: sales increase by 16.3% and conversion rate rises by 21.7% (both significant at the 1% level). Using Human Reply as the control (Column 2), the GenAI Reply treatment shows no statistically significant differences in either sales or conversion, suggesting that the GenAI chatbot matches the quality of human service but does not outperform it. When the No Service control is compared with the GenAI+Human Reply treatment (Column 3), the gains are even larger—sales improve by 25.0% and conversion by 29.0%—indicating complementarities between GenAI and human agents. By contrast, relative to the Human Reply control (Column 4), the GenAI+Human Reply treatment yields a marginally significant 11.5% increase in sales, with no statistically significant change in conversion (4.8%), implying that the hybrid approach can enhance revenue even when benchmarked against human agents. Finally, Table C2 shows that all estimated effects on cart value are statistically insignificant.
Table C1: Effects of Pre-sale Service Chatbot on Sales Across Experiments
@@
%Change 16.3% 3.7% 25.0% 11.5%
Observations 44,614 44,736 30,345 30,467
R-squared 0.000 0.000 0.000 0.000
  • 1
    “Sales” represents the total expenditure on product orders.
  • 2
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Table C2: Conversion Rate and Cart Value in Pre-sale Service Chatbot Experiments
Control: No Service Human Reply No Service Human Reply
(1) (2) (3) (4)
Extensive margin: Conversion Rate
Treat 0.0131*** -0.000768 0.0175*** 0.00358
(0.00256) (0.00261) (0.00295) (0.00301)
%Change 21.7% -1.0% 29.0% 4.8%
Observations 44,614 44,736 30,345 30,467
R-squared 0.001 0.000 0.001 0.000
Intensive margin: Cart Value (conditional on purchase)
Treat -1.264 1.220 -0.859 1.624
(1.036) (0.929) (1.203) (1.078)
%Change -4.5% 4.8% -3.1% 6.4%
Observations 3,076 3,300 2,092 2,316
@@
    The first block reports the Conversion Rate (binary indicator equal to 1 if consumer makes at least one order during the experimental period). The second block reports Cart Value, expenditure per consumer conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement
@@
The results are summarized in Table C3. Column 1 indicated no significant differences in product views between the two groups. However, treatment group consumers generated 1.10% more clicks (Column 2), spent 2.93% more (Column 4), and were 1.15% more likely to make a purchase (Column 5). We find no significant impact on cart value and purchase frequency (Columns 6 and 7).
We also explore further mechanism analysis of the Search Query Refinement application. In Table C4, we find that the likelihood of a product click increases by 0.3% ($p<0.01$; Column 1), consistent with improved search performance in combination with the documented purchase conversion increase in our main effect. Columns 2 and 3 further indicate that treated consumers view fewer products prior to clicking or purchasing, suggesting a reduction in search intensity. In addition, we observe a statistically significant 2.0% increase in the click-through rate ($p<0.01$; Column 4), defined as the ratio of product clicks to product views, suggesting that consumers found the exposed products more appealing and chose to seek additional details after viewing the summarized search results. The click-to-order conversion rate (Column 5), defined as the ratio of orders to clicks, remains statistically insignificant. This pattern echoes the fact that query refinement influenced only the composition of products retrieved immediately after a query search, not the information displayed on product detail pages. Overall, GenAI-facilitated query refinement enhanced the search algorithm’s ability to satisfy consumers’ demand, resulting in more effective matching and improved consumer purchases.
@@
(1) (2) (3) (4) (5) (6) (7)
Conversion Cart Purchase
Views Clicks Orders Sales Rate Value Frequency
Treat -0.549 0.0901*** 0.00154 0.0648** 0.00101** 0.370 -0.00315
(0.887) (0.0245) (0.00107) (0.0314) (0.000411) (0.334) (0.00842)
%Change -0.18% 1.10% 0.94% 2.93% 1.15% 1.47% -0.17%
Observations 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 163,381 163,381
R-squared 0.038 0.041 0.019 0.004 0.030 0.010 0.004
@@
    “Views” stands for the number of product views in the summarized search results pages. “Clicks” stands for the number of product clicks into product detail pages. “Orders” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Purchase Frequency” refers to the number of orders per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Table C4: More Exploration of Search Query Refinement
(1) (2) (3) (4) (5)
Views Views Click- Click-to-Order
Conditional on Conditional through Conversion
Is Click Click Order Rate Rate
Treat 0.00267*** -1.728* -10.34* 0.000767*** -0.000088
(0.00057) (1.034) (5.362) (0.000079) (0.000182)
%Change 0.33% -0.46% 1.48% 2.02% -0.35%
Observations 1,849,382 1,508,873 163,381 1,849,382 1,508,873
R-squared 0.013 0.039 0.069 0.003 0.014
  • 1
    “Is Click” measures consumers’ likelihood of making a click. It is a binary indicator for click, which equals 1 if a consumer makes at least one click during the experimental period, and 0 otherwise. “Views Conditional on Click” refers to the product views per consumer, conditional on making a click. “Views Conditional on Order” refers to the product views per consumer, conditional on making a purchase. “Click-through Rate” is the ratio of the number of product clicks to the number of product views, measuring the degree of conversion from views to clicks. “Click-to-Order Conversion Rate” stands for the ratio of the number of product orders to the number of product clicks, measuring the degree of conversion from clicks to purchases.
  • 2
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Product Description
@@
Column 1 of Table C5 shows no statistically significant differences in product clicks between the control and treatment groups. However, conditional on consumers clicking through to product detail pages—where the descriptions are displayed—treated consumers place 1.08% more orders and spend 2.05% more on those orders (Columns 2 and 3). This improvement is also reflected in a 1.27% increase in the conversion rate (Column 4), indicating a higher likelihood of purchase following exposure to AI-generated descriptions. By contrast, we detect no statistically significant changes in cart value or purchase frequency. Taken together, these results suggest that augmenting human-generated product descriptions with AI-generated content primarily affects purchase incidence rather than spending intensity, leading to higher overall sales.
Moreover, we stratify products based on the length of their original, human-generated descriptions, distinguishing between those with no or insufficient text (fewer than 50 words) and those with sufficiently detailed descriptions (more than 50 words). Products in the former group experience a 6.1% increase in sales following augmentation with AI-generated content ($p<0.05$), whereas products in the latter group show no significant effect (Columns 1 and 2 of Table C6), indicating that richer descriptions facilitate consumer decision-making. [^25] Based on internal research and expert surveys conducted by our partner platform, descriptions containing fewer than 50 words are classified as providing insufficient textual information. [^26] The analysis is restricted to English-speaking consumers, as description-length data are available only for English-language content.